### v1

# R1-Zero’s “Aha Moment” in Visual Reasoning on a 2B Non-SFT Model

Abstract
Recently DeepSeek R1 demonstrated how reinforcement learning with simple rule-based incentives can enable autonomous development of complex reasoning in large language models, characterized by the "aha moment", in which the model manifest self-reflection and increased response length during training. However, attempts to extend this success to multimodal reasoning often failed to reproduce these key characteristics. In this report, we present the first successful replication of these emergent characteristics for multimodal reasoning on only a non-SFT 2B model. Starting with Qwen2-VL-2B and applying reinforcement learning directly on the SAT dataset, our model achieves 59.47% accuracy on CVBench, outperforming the base model by approximately ~30% and exceeding both SFT setting by ~2%. In addition, we share our failed attempts and insights in attempting to achieve R1-like reasoning using RL with instruct models. aiming to shed light on the challenges involved. Our key observations include: (1) applying RL on instruct model often results in trivial reasoning trajectories, and (2) naive length reward are ineffective in eliciting reasoning capabilities. The project code is available at https://github.com/turningpoint-ai/VisualThinker-R1-Zero
Work in Progress This research is still a work in progress. In this report, we share our successes, failures, and insights from our preliminary exploration of potential approaches for developing R1-like reasoning in multimodal models. This is an ongoing effort and we will be committed to continuous update and improvement. This report presents only a minimal version, with the full version to be released soon.
1 Introduction
Figure 1: The training dynamics of VisualThinker-R1-Zero on Qwen2-VL-2B base model. Benchmark accuracy is measured on CV-Bench, and the average response length is calculated from rollouts on SAT training samples. Initially, we observed a drop in length because the base model tended to generate HTML code. This behavior was quickly suppressed by RL, leading the model to adopt a more appropriate output format and a regular increase in response length. Afterwards, we observed a multimodal ‘aha moment’—the emergence of self-reflection in models’ response, as described in the DeepSeek-R1 paper, followed by a consistent positive correlation between response length and benchmark accuracy.
Recently, DeepSeek R1 [2] has demonstrated how Reinforcement Learning (RL) with simple rule-based incentives can enable a large language model to build complex reasoning capabilities autonomously. A key finding from this work was the emergence of advanced reasoning patterns without explicit supervision—what the researchers termed the "aha moment," characterized by self-reflection, and spontaneous increase in response length during training as the model learned to explore increasingly sophisticated problem-solving strategies.
Many researchers [1; 20; 13; 17] have attempted to extend this success to multimodal reasoning, where models process and reason on both visual and textual information. However, these efforts have primarily struggled to reproduce the key characteristics exhibited by DeepSeek R1 mentioned above—specifically, the emergent "aha moment" and increased response length during reasoning. These implementations often failed to demonstrate the autonomous development of sophisticated reasoning strategies observed in DeepSeek R1’s training.
In this report, we present the first successful replication of these key characteristics for multimodal reasoning on only a non-SFT 2B model. With both "aha moment" and increased length (Figure 1), our approach demonstrates that direct application of reinforcement learning on non-sft model can induce sophisticated reasoning capabilities even in smaller multimodal models without supervised fine-tuning. We start from the Qwen2-VL-2B [16] base model and directly perform reinforcement learning. Without any SFT, our model achieves 59.47% accuracy on CVBench [15], outperforming the base model by approximately ~30% and exceeding the SFT model by ~2%.
In addition to our replication on non-sft model, we also share our insights and failed attempts in achieving R1-like reasoning using RL with instruct model. We observed that starting from a supervised fine-tuned model often failed to reproduce the observations and findings reported by DeepSeek-R1. Upon investigation on this issue, we found that (1) Despite improved performance, RL on instruct model leads to superficial reasoning rather than genuine problem-solving strategies, and (2) naive length reward are ineffective at inducing deeper reasoning capabilities.
In summary, our key contributions are as follows:
  • We are the first team to replicate the key characteristics of R1 success ("aha moment" and increased reasoning length) on multimodal reasoning tasks with a non-SFT 2B model.
  • We showed that vision-centric spatial reasoning tasks could also benefit from improved reasoning capabilities.
  • We demonstrated that applying RL on instruction-tuned models leads to superficial reasoning.
  • We open-sourced our project to facilitate future studies on multimodal reasoning.
2 Related Works
2.1 Multimodal Reasoning
Researchers have demonstrated that LLMs can be post-trained to elicit enhanced reasoning abilities [8]. With pre-trained visual encoders to understand visual content alongside textual data, multimodal LLM’s reasoning abilities are investigated and enhancement typically requires sophisticated prompting designs [6] or large amounts of reasoning training data [14; 18]. The research community is increasingly interested on developing more natural methods to incentivize higher intelligence in models without relying on extensive supervised data or complex prompting techniques.
2.2 The "Aha Moment" Phenomenon in DeepSeek R1
A recent breakthrough study, DeepSeek R1 [2], demonstrated that reinforcement learning can incentivize a model’s reasoning abilities without any supervised reasoning data. Intriguingly, researchers discovered an "aha moment" when directly applying RL with rule-based reward on mathematical datasets—the model autonomously developed advanced problem-solving strategies, including reflection and self-correction.
We summarize the key characteristics that contributed to DeepSeek R1’s success and compare them with our model and other multimodal replications in Table 1. Specifically, we highlight two emergent phenomena: the "aha moment" and increasing response length. The "aha moment" refers to the model’s autonomous development of advanced problem-solving strategies during training, while the increasing response length indicates the model naturally learns to allocate more thinking time for reasoning tasks. It remains questionable whether existing multimodal replications [4; 1; 3] without these key characteristics can be considered truly valid implementations of the R1 approach.
Table 1: Comparison between DeepSeek R1 and its multimodal replications
Feature DeepSeek R1 [2] VisualThinker R1 Zero (Ours) R1-V [1] R1-Multimodal-Journey [4] open-r1-multimodal [3]
Base Model DeepSeek V3 Qwen2-VL-2B Qwen2-VL-2B-Instruct Qwen2-VL-2B-Instruct Qwen2-VL-2B/7B-Instruct
Modality Language Vision + Language Vision + Language Vision + Language Vision + Language
Aha Moment Yes Yes No Yes No
Response Length Dynamics $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
3 VisualThinker R1 Zero
In this part, we demonstrate VisualThinker-R1-Zero have showed an emergence of “aha moment” by directly applying RL training to the non-SFT base model, leading to a superior performance on vision-centric tasks on non-SFT 2B models.
Base Model
Our method builds on Qwen-2-VL-2B [16] as the base model, applying GRPO [12] with a tailored chat template and prompting strategy to enhance its reasoning capabilities. We posit that applying GRPO to the base model may be a more efficient and effective way to replicate multimodal R1 reasoning compared to training instruct fine-tuned model. As we will show in Section 5, the instruction-tuned variant, Qwen-2-VL-2B-Instruct, demands significantly more training resources as a base model yet struggles to replicate the emergent "aha moment" observed in DeepSeek R1 with notable failure modes.
Training Recipe
We train the model directly on the SAT [10] dataset to let the base model to explore various spatial reasoning for each question $q$ in the dataset $Q$, we apply the following chat template:
[…]
For each of the question $q$ in the dataset, the model generate an response $o$ with this prompt template and it is then optimized using RL objective.
RL algorithm
Existing repos applying RL on top of fine-tuned visual models failed to replicate DeepSeek r1’s key characteristics. In contrast, we witness prolonged reasoning trajectory and “aha moment” with an overlooked approach that directly applies GRPO [12] to an non-SFT model. Our findings suggest that this setting is the key to TRUE “aha moment” in multimodal reasoning. Now we briefly review the GRPO algorithm we adopted for RL training.
To ease the burden of training an additional value function approximation model employed by PPO [11], GRPO adopted the average reward of sampled response of the policy model as the baseline in computing the advantage. Specifically, given the an input question $q$, we first sample a group of responses $\{o_{1},o_{2},\cdots,o_{G}\}$ and compute corresponding rewards $\{r_{1},r_{2},\cdots,r_{G}\}$ with the reward model. The advantage is then computed as:
$\hat{A}_{i,t}=\widetilde{r}_{i}=\frac{r_{i}-{\rm mean}(\mathbf{r})}{{\rm std}(\mathbf{r})}$ (1)
The policy model is then optimized by maximizing the following KL objective:
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\\ &\quad\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\Bigg{]}-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\Bigg{\}}\Bigg{]}\end{split}$ (2)
where $\pi_{\theta}$ and $\pi_{old}$ are the current and old policy, and $\epsilon$ and $\beta$ are hyper-parameters introduced in PPO.
Reward Modeling
Following DeepSeek-R1, our RL approach remains elegant, avoiding the use of reward models [9] or Monte Carlo tree search (MCTS)-like techniques [19]. Specifically, we employ a rule-based reward function that evaluates responses based on their format and correctness:
  • If the response provides a final answer and is correct, the model receives an accuracy reward of +1.
  • If the response encloses its thinking in <think></think> and the final answer in <answer></answer> tags, the model receives a format reward of +1.
  • otherwise, the model receives 0 reward.
Our implementation is based on DeepSeek R1’s report. Initial experiments suggest that this reward function helps the policy model quickly converge towards generating responses in the desired format.
4 Experiments
In this study, we demonstrate that small non-sft model could solve vision-centric spatial reasoning task with our method, a type of benchmark often pose challenges to even larger models. We train our models on SAT [10], a VQA dataset comprising 218k question-answer pairs synthesized using a photo-realistic physics engine to enhance spatial intelligence. Our training focuses on the static subset, which includes questions on relative spatial relationships, relative depth, and object counting.
4.1 Evaluation
To test the generalization of our method, we evaluate on CVBench [15], a realistic vision-centric benchmark consists of 2,638 examples repurposed from standard vision datasets on fundamental 2D and 3D reasoning tasks. CVBench formulates natural language questions to assess spatial relationships, object counting, depth ordering, and relative distance. This setup allows us to systematically examine how well our recipe could improve spatial reasoning capability.
4.2 Implementation Details
We conduct all our experiments using four NVIDIA H100 GPUs (80GB each), setting the batch size to 1 per device. The model is trained for 1500 steps with a learning rate of $1\times 10^{-6}$ and a temperature of 1.0. We found long response length is a must for observing increasing response length during training, so we set the maximum response length as 700. During GRPO optimization, we sample 8 responses per step and apply a KL coefficient of 0.04.
Table 2: Hyper-parameters of VisualThinker R1 zero GRPO training.
Setting Value
Batch Size per Device 1
Gradient Accumulation Steps 2
Training Steps 1500
Learning Rate $1\times 10^{-6}$
Temperature 1.0
Maximum Response Length 700
Number of Responses per GRPO Step 8
KL Coefficient 0.04
To demonstrate the superior effectiveness of directly applying GRPO on base model, we compare our method with the non-sft pre-trained mode as the baseline, and compare our method against the base model SFT on the same SAT dataset.
4.3 Main Results
Benchmark Results
In the experiments, we fine-tuned the Qwen2-VL-2B non-sft model and evaluated its performance on CV-Bench. During training, we can observe the model autonomously develop increasing response length and the performance comes with it in Figure 1 on CV-Bench. We can also observe that directly applying RL on base model has superior performance compared to SFT method in Figure 2.
In addition, we tested our method on various spatial reasoning dataset including BLINK [5] [^1] Following SAT [10] evaluation, we use two spatial splits of BLINK - Multiview reasoning, Relative Depth, and Spatial Relations. and VSR [7], illustrated in Table 3. Our method demonstrated improved performance over the Qwen2-VL-2B (base) by ~30%, and the Qwen2-VL-2B SFT(base + SFT) by ~2% on CV-Bench. We also achieve superior performance on the BLINK and VSR benchmark, with our method achieves around 27% advantage comparing against the model trained with SFT. This suggests that visual reasoning could significantly benefit from R1-Zero training, demonstrating more scalable training through RL’s exploration of diverse reasoning.
Figure 2: Comparison between RL and SFT training. Our method achieves a significant improvement over the base model and the instruction fine-tuned model. Specifically, Qwen2-VL-2B + R1 outperforms Qwen2-VL-2B (base model) by approximately ~30%, Qwen2-VL-2B-Instruct (instruction fine-tuned model) by ~5%, and Qwen2-VL-2B SFT (base model + SFT) by ~2%.
Table 3: Results on vision-centric benchmarks. Table shows RL training on base model has overall better performance over SFT training and the base model.
Model CV-Bench BLINK VSR
Count Acc(%) Relation Acc(%) Depth Acc(%) Distance Acc(%) Total Acc(%) Relative Depth Acc(%) Spatial Relation Acc(%) Average Acc(%) Average Acc(%)
Qwen2-VL-2B 54.69 22.46 0.16 31.66 31.38 13.70 0.69 6.74 0.0
Qwen2-VL-2B + SFT 60.02 68.92 55.00 45.83 57.84 58.06 47.55 52.43 35.80
Qwen2-VL-2B + GRPO (Ours) 59.64 66.76 54.16 56.66 59.47 50.80 55.94 53.18 62.32
Multimodal Aha Moment
A particularly intriguing phenomenon observed during the training of DeepSeek-R1-Zero is the occurrence of an “aha moment”:
**Wait, wait. Wait. That’s an aha moment I can flag here.**
Let’s reevaluate this step-by-step to identify if the correct sum can be...
We started with the equation:
This aha moment indicates DeepSeek-R1-Zero spontaneously builds a reasoning strategy rethinking its initial approach for improved reasoning capability.
Beyond strong performance gain, we also observed our model exhibit very interesting "aha moment" aligned with the finding in DeepSeek R1: during training, the model spontaneously revisit its previous judgment and explore alternative options:
Therefore, dark brown wooden bed with white blanket is not above
the doorway.
**But wait! I can think of something else.** <think>
5 Challenges of Applying RL to Supervised Fine-Tuned Models
Starting from scratch with a non-SFT multimodal model, we have demonstrated that RL empowers VisualThinker-R1-Zero to attain robust reasoning capabilities without any supervised fine-tuning data. In addition to this replication, one might be inclined to apply RL directly to supervised fine-tuned models given its stronger instruction following capability.
In this section, we share our failed cases on applying GRPO algorithm described in Section 3 to the Qwen2-VL-2B-Instruct model (a supervised fine-tuned Qwen VL model) to provide insights that may benefit future research. However, note that this does not imply that these approaches are incapable of building effective visual reasoning models. Specifically, we found that starting from a supervised fine-tuned model exposes problems of trival reasoning patterns. We then demonstrate our findings in (1) investigating the reason and implication of this phenomenon and (2) attempting to incentivize sophisticated reasoning.
5.1 Emergence of Trivial Reasoning Patterns
Applying RL on an SFT model does improve the model’s performance on CVBench (Table 4). However, it is questionable whether this method incentivizes the models to achieve higher intelligence, as Figure 3 demonstrates the degeneration of model responses into meaningless or trivial reasoning patterns. The observed reasoning trajectory follows a structure: a trivial and generic strategy within <think></think> tags followed by a question-specific answer between <answer></answer> tags. One hypothesis is that the model enhances its performance during RL training without necessarily developing genuine reasoning capabilities.
Trivial Reasoning Trajectory Question: Which object is closer to the camera taking this photo, the box (highlighted by a red box) or the desk (highlighted by a blue box)?
Response: <think>To determine which object is closer to the camera, I will measure the distance from the camera to each object.</think> <answer>box</answer>
Figure 3: Example response of applying RL to supervised fine-tuned models.
5.2 Preliminary Investigation of Trivial Reasoning Trajectories
Given the trivial reasoning trajectories, a nature question is whether performance improvements occur through enhancement of the vision encoder during training. We hypothesized that when we freeze the vision encoder during RL, the model may focus more on developing sophisticated reasoning strategies; conversely, when we freeze the LLM during RL, the performance would be the same level as the full fine-tuning setting. Surprisingly, results in Table 5 demonstrate that both approaches achieve greater improvement than the vanilla implementation yet they still generate short and trivial responses. This observation suggests the complexity of training dynamics of RL on multimodal models and further analysis is needed to better understand the phenomenon.
Table 4: Evaluation of applying RL to instruct models. Prompting with only <answer> achieves superior performance in the context of applying RL on instruct models.
Prompting Strategy Total Acc (%) Count Acc (%) Relation Acc (%) Depth Acc (%) Distance Acc (%)
Instruct Model - 55.64 45.43 68.92 58.66 51.66
Instruct Model + RL with reason 66.03 69.54 61.84 66.50 65.50
Table 5: Evaluation of applying RL to instruct models with different freezing components. Either freezing language or vision components of instruct model can both leads to performance improvement.
Freezing Components Total Acc (%) Count Acc (%) Relation Acc (%) Depth Acc (%) Distance Acc (%)
Instruct model + RL - 63.57 61.68 71.08 61.50 60.00
Instruct model + RL vision encoder 68.34 64.21 78.00 65.67 66.00
Instruct model + RL LLM 65.35 62.05 72.46 65.17 62.17
Figure 4: Response length across training steps for different fine-tuning settings during RL. The x-axis represents training steps, while the y-axis shows the response length. Models with different fine-tuning configurations are compared: Freeze LLM (green), Freeze Vision Encoder (blue), and Full Finetune (red). The response length drops significantly in the early training phase and stabilizes over time. However, despite improved accuracy, all three RL-based fine-tuning on Instruct Model does not necessarily enhance reasoning capabilities, as responses tend to remain short and trivial
5.3 Failed Attempts with Length Reward
The key characteristics of DeepSeek R1’s success is its increased response length during training. After observing our failure to replicate DeepSeek R1’s reasoning capabilities when starting from instruction-tuned models, we investigated whether we could encourage more sophisticated reasoning patterns by directly rewarding longer response.
Figure 5: Model performance during training with length-rewarded RL on instruction ginetuned models. Despite implementing length rewards to encourage longer responses, model performance shows no improvement in reasoning quality as response length increases. While average response length (red line) grows substantially, benchmark accuracy (green line) remains relatively stable, suggesting that longer responses do not necessarily translate to enhanced reasoning capabilities.
Length Reward Hacking Example Question: Considering the relative positions of the giraffe (annotated by the red box) and the person in the image provided, where is the giraffe (annotated by the red box) located with respect to the person?
Model Response: <think> - The giraffe (annotated by the red box) is located in the foreground of the image. - The person is in the background, standing behind the giraffe. - The giraffe is closer to the viewer than the person. - The person is not directly above the giraffe. - The person is not directly below the giraffe. - The giraffe is not above the person. - The giraffe is not below the person. - The person is not in front of the giraffe. - … [[[[Repeats until it reaches max_new_tokens limit]]]] … </think>
Figure 6: Response example of model hacking naive length reward. Model hacks reward by repeating trivial response.
Thus, we implemented a length-based reward mechanism: in addition to the vanilla accuracy and format rewards described in Section 3, we added an auxiliary reward of $+0.001$ for each additional token generated. However, as shown in Figure 5, naively rewarding lengthy responses does not improve model performance and often leads to reward hacking behaviors, with models generating extremely long yet meaningless content as illustrated in Figure 6. This finding suggests that the emergent reasoning capabilities observed in DeepSeek R1 cannot be easily achieved through enforcing long response length during the training of instruction-tuned models.
6 Conclusion
This paper presents VisualThinker R1 Zero, the first successful multimodal replication of DeepSeek R1’s emergent reasoning characteristics. By applying reinforcement learning directly to a non-fine-tuned Qwen2-VL-2B model, we observed both the "aha moment" and increased response length during training—key indicators of autonomous reasoning development. Empirically, our approach achieved 59.47% accuracy on CVBench, outperforming both base and instruction-tuned models without any supervised fine-tuning. We also share our insights and failed attempts in achieving R1-like reasoning using RL with instruct model: applying RL to supervised fine-tuned models leads to trivial reasoning trajectories rather than genuine problem-solving strategies.
This report is a work in progress for presenting our preliminary findings, we plan to release further updates with deeper investigations and insights to continuously expand and refine this report with our ongoing exploration of the technical roadmap for realizing R1-like multimodal reasoning.